Exercise 5.56 - Enhanced - with Feedback and Hints A sample of N20 gas has a density of 2.65 g L-at 298 K. Part A What is the circle answers
Answer #1

Molar mass of N2O,

MM = 2*MM(N) + 1*MM(O)

= 2*14.01 + 1*16.0

= 44.02 g/mol

Lets derive the equation to be used

use:

p*V=n*R*T

p*V=(mass/molar mass)*R*T

p*molar mass=(mass/V)*R*T

p*molar mass=density*R*T

P*44.01 = 2.65*0.08206*298

P*44.01 = 64.80

P = 1.47 atm

Answer: 1.47 atm